The population mean waiting time of customers of the XYZ Burger chain is 3.4 minutes with a population standard deviation of 1.2 minutes. The customer service department found in a sample of 50 customers that the mean waiting time was 3.2 minutes. At the 4% significance level, can we conclude that the mean waiting time is different from the population mean?
Answer by stanbon(57290)

The population mean waiting time of customers of the XYZ Burger chain is 3.4 minutes with a population standard deviation of 1.2 minutes. The customer service department found in a sample of 50 customers that the mean waiting time was 3.2 minutes. At the 4% significance level, can we conclude that the mean waiting time is different from the population mean?
---------
Ho: u = 3.4
Ha: u is not 3.4
---------------------
x-bar = 3.2
---
t(3.2) = (3.2-3.4)/[1.2/sqrt(50)] = -1.1785
----
p-value = 2*P(t< -1.1785 when df = 49) = 0.2443
---
Since the p-value is greater than 4%, fail to reject Ho.
The test results support the Ho claim at the 4% significance level.
Cheers,
Stan H.